OpenStudy (anonymous):

Find equations of both lines through the point (2,-3) that are tangent to the parabola y+x^2+x. Help would be much appreciated :)

OpenStudy (anonymous):

oops y=x^2+x

OpenStudy (anonymous):

i hope to god you can use calculus

OpenStudy (anonymous):

you can do it without, but it is a real pain

OpenStudy (anonymous):

yeah, no prob!

OpenStudy (anonymous):

ok whew

OpenStudy (anonymous):

so the derivative is \(2x+1\) and if the point is tangent to the graph, that means that the slope has to agree with \(2x+1\) for whatever \(x\) value you have the slope through a point on the graph \(y=x^2+x\) looks like \((x,x^2+x)\) and the slope between that point and \((2,-3)\) will be \[\frac{x^2+x+3}{x-2}\]

OpenStudy (anonymous):

ok i said that wrong, but you get the idea i hope in any case, your job is to solve \[\frac{x^2+x+3}{x-2}=2x+1\] for \(x\) you can see that you will get a quadratic equation, which is why you have two solutions

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

yw sorry my explanation was so convoluted, hope it makes sense

OpenStudy (anonymous):

it did :)

OpenStudy (anonymous):

Here is solution not using Calculus Let g(x)= -3 + m(x-2) be any line going through the point (2,-3) Let us find the points of intersections of this line with the parabola. We have to slove \[ x^2 + x = -3 + m(x-2)\\ x^2 +x(1-m)+3+2m==0\\ \] to be tangent, the two roots have to be equal. Hence the discriminant should be zero \[ (1-m)^2 -4(3+2m)==0\\ m=-1\\ m=11 \] This will give you two tangents \[ y1=-1-x\\ y2=-3+11 (-2+x)\\ \]

OpenStudy (anonymous):

@satellite73